# MetaSym: A Symplectic Meta-learning Framework for Physical Intelligence

### v1

@@ B.3 Quadrotor
The quadrotor system challenges data-driven methods with its floating-base dynamics. To generate the training and OOD validation datasets we use Crocoddyl trajectory optimization package based on the dynamics model proposed by (Geisert & Mansard, 2016). The training dataset comprises 30 systems with randomized parameters such as inertia, torque constant and rotor lengths as indicated in Table 7. Each trajectory is generated from a randomized initial condition to a random terminal position with zero velocity, within a pre-set bounding box, to avoid unrealistic velocities.
In the same manner the OOD validation set contains 10 trajectories, each one corresponding to a system with parameters drawn from the distributions indicated in Table 7.
Parameters Values
Inertia $\mathcal{I}_{diag}$ $\mathrm{Uniform}(0.1,0.2)$
mass (m) $\mathrm{Uniform}(0.5,3.0)$
$d_{cogs}$ $\mathrm{Uniform}(0.2,0.5)$
$C_{torques}$ $\mathrm{Uniform}(0.001,0.1)$
Table 6: In-distribution parameters $dt=0.01$, $T=250$.
Parameters Values
Inertia $\mathcal{I}_{diag}$ $\mathrm{Uniform}(0.3,0.7)$
mass (m) $\mathrm{Uniform}(3.0,5.0)$
$d_{cogs}$ $\mathrm{Uniform}(0.5,0.7)$
$C_{torques}$ $\mathrm{Uniform}(0.1,0.2)$
Table 7: Out-of-distribution parameters $dt=0.01$, $T=500$
Appendix C Near Symplectic Architecture
In this section, we prove that composing our SymplecticEncoder $\Phi_{\theta_{SE}}$ with our ActiveDecoder $\Phi_{\theta_{AD}}$ handling control signals and dissipative effects yields a near-symplectic map, with an explicit bound on preserving approximate symplectic geometry. This formalizes the realistic effects we can introduce to MetaSym and the inductive bias provided by symplectic invariants.
@@ C.1 Overall Map
We aim to show that if $\Phi_{\theta_{AD}}$ remains a small (bounded) perturbation from identity in $\mathbf{z}_{c}$-space, then $\Phi_{\theta}$ preserves the symplectic structure up to a small error term. This is called near-symplecticity. We assume that the dissipation $\mathbf{d}_{t}$ is seperable for the sake of this proof.
@@ C.2 SymplecticEncoder
From Jin et al., we know that LA-SympNets are fully symplectic due to their construction (Jin et al., 2020). By extension, we can show that our SymplecticEncoder $\Phi_{{\theta}_{SE}}$.
@@ C.3 Decoder with Control and Dissipation
C.4 Near-Symplectic Proof and Explicit Bound
In this section, we want to prove that if $\mathbf{z}_{c}\mapsto\mathbf{z}_{c}+F_{\theta_{AD}}(\mathbf{z}_{c})$ is a bounded perturbation from identity (in partial derivatives), then the composition with a strict symplectic map remains close to preserving $J$.
@@ C.4 Near-Symplectic Proof and Explicit Bound
For the ActiveDecoder map, we assume the partial derivative w.r.t. $\mathbf{z}_{c}$ in $F_{\theta_{AD}}$ is bounded by a linear-type function of $(\|\mathbf{u}_{t}\|,\|\mathbf{d}_{t}\|)$:
@@
$\mathrm{d}\Phi_{\theta}(\mathbf{x},\mathbf{u}_{t},\mathbf{d}_{t})=\;\underbrace{\mathrm{d}_{\mathbf{z}_{c}}\!\Phi_{\theta_{AD}}\bigl{(}\mathbf{z}_{c},\mathbf{u}_{t},\mathbf{d}_{t}\bigr{)}}_{\textstyle I+A,\;\|A\|\leq\rho}\;\;\times\;\;\underbrace{\mathrm{d}\Phi_{\theta_{SE}}(\mathbf{x})}_{\textstyle\text{strictly symplectic.}}$
Theorem C.2 (Near-Symplectic Composition).
@@
giving the same $\rho=\alpha_{0}+\alpha_{u}U_{\max}+\alpha_{d}D_{\max}$. Provided $\rho\ll 1$, $\Phi_{\theta}$ remains near-symplectic.

### v2

@@ B.3 Quadrotor
The quadrotor system challenges data-driven methods with its floating-base dynamics. To generate the training and OOD validation datasets we use Crocoddyl trajectory optimization package based on the dynamics model proposed by [52]. The training dataset comprises 30 systems with randomized parameters such as inertia, torque constant and rotor lengths as indicated in Table 4(a). Each trajectory is generated from a randomized initial condition to a random terminal position with zero velocity, within a pre-set bounding box, to avoid unrealistic velocities.
In the same manner the OOD validation set contains 10 trajectories, each one corresponding to a system with parameters drawn from the distributions indicated in Table 4(b).
Parameter Value
Inertia $\mathcal{I}_{\text{diag}}$ $\mathrm{Uniform}(0.1,0.2)$
mass $m$ $\mathrm{Uniform}(0.5,3.0)$
$d_{\text{cogs}}$ $\mathrm{Uniform}(0.2,0.5)$
$C_{\text{torques}}$ $\mathrm{Uniform}(0.001,0.1)$
(a) In-distribution parameters. Time step $dt=0.01$, duration $T=250$.
Parameter Value
Inertia $\mathcal{I}_{\text{diag}}$ $\mathrm{Uniform}(0.3,0.7)$
mass $m$ $\mathrm{Uniform}(3.0,5.0)$
$d_{\text{cogs}}$ $\mathrm{Uniform}(0.5,0.7)$
$C_{\text{torques}}$ $\mathrm{Uniform}(0.1,0.2)$
(b) Out-of-distribution parameters. Time step $dt=0.01$, duration $T=500$.
Table 4: Parameter distributions for simulating quadrotor dynamics. The in-distribution set covers training conditions, while the out-of-distribution set reflects variations in inertia, mass, center of gravity shift, and torque coupling beyond the training domain.
Appendix C $O(\rho)$ Near-Symplectic Architecture
In this section, we prove that composing our SymplecticEncoder $\Phi_{\theta_{SE}}$ with our ActiveDecoder $\Phi_{\theta_{AD}}$ handling control signals and dissipative effects yields a near-symplectic map, with an explicit bound on preserving approximate symplectic geometry even when tested in adverse dissipative conditions. This formalizes the realistic effects we can introduce to MetaSym and the inductive bias provided by symplectic invariants.
@@ C.1 Overall Map
We aim to show that if $\Phi_{\theta_{AD}}$ remains a small (bounded) perturbation $O(\rho)$ from identity in $\mathbf{z}_{c}$-space, then $\Phi_{\theta}$ preserves the symplectic structure up to a small error term. This is called $O(\rho)$ near-symplecticity. We assume that the dissipation $\mathbf{d}_{t}$ and control-input $\mathbf{u}_{t}$ are seperable for the sake of this proof.
@@ C.2 SymplecticEncoder
From Jin et al., we know that LA-SympNets are fully symplectic due to their construction [12]. By extension, we can show that our SymplecticEncoder $\Phi_{{\theta}_{SE}}$ is symplectic.
@@ C.3 Decoder with Control and Dissipation
C.4 $O(\rho)$ Near-Symplectic Proof and Explicit Bound
In this section, we prove that if $\mathbf{z}_{c}\mapsto\mathbf{z}_{c}+F_{\theta_{AD}}(\mathbf{z}_{c})$ is a bounded perturbation from identity (in partial derivatives), then the composition with a strict symplectic map remains close to preserving $J$.
@@ C.4 $O(\rho)$ Near-Symplectic Proof and Explicit Bound
For the ActiveDecoder map, we take the partial derivative w.r.t. $\mathbf{z}_{c}$ in $F_{\theta_{AD}}$ is bounded by a linear-type function of $(\|\mathbf{u}_{t}\|,\|\mathbf{d}_{t}\|)$:
@@
$\mathrm{d}\Phi_{\theta}(\mathbf{x},\mathbf{u}_{t},\mathbf{d}_{t})=\underbrace{\mathrm{d}_{\mathbf{z}_{c}}\!\Phi_{\theta_{AD}}\bigl{(}\mathbf{z}_{c},\mathbf{u}_{t},\mathbf{d}_{t}\bigr{)}}_{\textstyle I+A,\;\|A\|\leq\rho}\quad\times\underbrace{\mathrm{d}\Phi_{\theta_{SE}}(\mathbf{x})}_{\textstyle\text{strictly symplectic.}}$
Theorem C.2 ($O(\rho)$ Near-Symplectic Composition).
@@
giving the same $\rho=\alpha_{0}+\alpha_{u}U_{\max}+\alpha_{d}D_{\max}$.
We choose not to explicitly bound this perturbation during training to allow for MetaSym to model extremely dissipative systems. In Appendix D.2, we provide the perturbation bound $C\rho$ emperically for systems undergoing extreme dissipation and control. We observe that this bound is $<1$ even under these extreme settings.
Appendix D Ablation Studies
To rigorously assess the efficacy of our proposed framework relative to current state-of-the-art (SoTA) methods, we conduct a series of ablation studies. All experiments are performed on a moderately complex yet tractable spring-mesh system consisting of a $3\times 3$ node grid. Unless otherwise noted, all other hyperparameters remain identical to those in Tables 2(a) & 2(b).
D.1 Modeling Conservative Systems
Our architecture is designed to explicitly preserve symplectic structure through a dedicated SymplecticEncoder, which captures the conservative phase-space dynamics. By design, it conserves invariant quantities such as energy and volume in phase space. The ActiveDecoder captures dissipative and control-related effects, modeled as bounded perturbations to the symplectic manifold (see Section C.4). These, effects however should not prevent the ActiveDecoder to adapt to conservative systems. To highlight this, using the pre-trained ActiveDecoder on the dissipative dataset with active dynamics, we test its accuracy on purely conservative spring-mesh systems. The achieved trajectory MSE across 25 conservative springmesh systems is $\mathbf{0.050\pm 0.016}$, showcasing our method’s adaptability and accuracy to conservative systems that may have chaotic behaviour.
D.2 Empirical Perturbation Bound Estimation
The ActiveDecoder models the dissipative and control-input effects of a system as perturbations to the symplectic manifold that the conservative dynamics of a given system evolve on. While Section C.1 provides a definitive argument in favor of this interpretation, calculating this bound analytically is not possible. To further support our claims we performed an ablation study that calculates this bound empirically, using the pretrained weights for a dissipative 3x3 spring-mesh system with parameters such as decay that are chosen to represent a real deformable surface such as polypropylene undergoing extreme deformation (eg. a trampoline). In this edge case, The bound $C\rho$ is calculated as below:
$\|\mathrm{d}\Phi_{\theta^{*}}^{\top}J\mathrm{d}\Phi_{\theta^{*}}-J\|_{2}=0.999$ (9)
where $J=\begin{pmatrix}0&I_{d}\\ -I_{d}&0\end{pmatrix}$ and $\Phi_{\theta^{*}}$ the trained network.
D.3 Effectiveness of Meta-Attention Mechanism
Our meta-learning framework adapts the Queries and Values of the cross-attention in the ActiveDecoder. Table 5 indicates the efficacy of our meta-attention framework compared to MetaSym without meta-attention across 3 different types of systems with realistic dissipation and inertial parameters. Note that we do not stress-test the model with as high-noise as in Section 4.4, since this ablation study focuses on the efficacy of our meta-learning method
Table 5: Meta-Learning comparison across different OOD systems.
3×3 Spring-Mesh Quantum System Quadrotor
MSE w/o Meta-Attention (±$\pm\sigma$) $0.613\;(0.432)$ 1.000 (0.200) 16.457 (15.184)
MSE with Meta-Attention (±$\pm\sigma$) $\boldsymbol{0.230\;(0.115)}$ 0.898 (0.241) 8.397 (5.906)
Finally, as showcased by Fig. 5 for the quadrotor, the fast-adaptation achieves smooth convergence as the overall training converges with significant loss minimization. This proves the smooth dynamics of our meta-learning framework and the absence of severe instabilities that could have hindered its performance.
Figure 5: Inner Adaptation convergence for early training stage (left) and close-to-convergence training stage (right) for a quadrotor.
Appendix E Compute Resources
The dataset for the 10x10 spring mesh system and the associated baselines and benchmarks were run on a Lambda cloud instance consisting of a NVIDIA A100 with 40GB of GPU memory and 200GB of RAM. All other experiments were run on a system containing 252GB of RAM, 32 core processor and a 24GB NVIDIA RTX 4090.
Appendix F Hyperparameters for Benchmarks
This section includes the hyperparameters of our framework, which we used to benchmark each system as outlined in Table 1. This ensures the reproducibility and completeness of our method.
Table 6: Hyperparameter settings for MetaSym when benchmarking the spring-mesh.
Hyperparameter Encoder Decoder
Optimizer
Outer: AdamW
Inner: Adam
Outer: AdamW
Inner: AdamW
Learning Rate
Outer: 0.001
Inner: 0.003
Outer: 0.007
Inner: 0.01
Inner Steps 3 10
Context Window N/A 30
Dropout 0.1* 0.1
Layers 3 1
Attention Heads N/A 4
Early Stopping Epochs 110 312
*DropConnect
Table 7: Hyperparameter settings for MetaSym when benchmarking the open-quantum system.
Hyperparameter Encoder Decoder
Optimizer
Outer: AdamW
Inner: Adam
Outer: AdamW
Inner: AdamW
Learning Rate
Outer: 0.001
Inner: 0.003
Outer: 0.008
Inner: 0.03
Inner Steps 3 15
Context Window N/A 30
Dropout 0.2* 0.2
Layers 3 1
Attention Heads N/A 4
Early Stopping Epochs 294 354
*DropConnect
Table 8: Hyperparameter settings for MetaSym when benchmarking the quadrotor.
Hyperparameter Encoder Decoder
Optimizer
Outer: AdamW
Inner: Adam
Outer: AdamW
Inner: AdamW
Learning Rate
Outer: 0.001
Inner: 0.003
Outer: 0.01
Inner: 0.008
Inner Steps 3 15
Context Window N/A 10
Dropout 0.4* 0.45
Layers 3 1
Attention Heads N/A 4
Early Stopping Epochs 34 222
*DropConnect